Two trains leave from the same point at the same time, traveling in opposite directions. One travels 13 mph faster than the other. After 4 hours, they are 484 mi apart. Find the rate of each train.
Answer by mathstutor494(120)

Let speed of train#1 be x mi/hr
So speed of train#2 would be (x+13) mi/hr
After 4 hours, they would have travelled 4x and 4(x+13) mi respectively
Given that after 4 hours, they are 484 mi apart, so
4x + 4(x+13)= 484
8x + 13 = 484
8x= 484-13 = 432
Hence x = 54 mi/hr
Ans : Rate (speed) of each train is 54 mi/hr and 54+13 = 67 mi/hr